# Quantifying non-Hermiticity using single- and many-particle quantum properties

### v1

@@ Acknowledgements.
This project is funded by the European Union under NextGenerationEU. PRIN 2022 Prot. n. 2022ATM8FY (CUP: E53D23002240006). This project has received funding from the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant R20PEX7Y3A). The project is funded under the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.4 - Call for tender No. 1031 of 17/06/2022 of Italian Ministry for University and Research funded by the European Union – NextGenerationEU (proj. nr. CN_00000013). This work was supported by the Swiss State Secretariat for Education, Research and innovation (SERI) under contract number UeMO19-5.1. Project DYNAMITE QUANTERA2_00056 funded by the Ministry of University and Research through the ERANET COFUND QuantERA II – 2021 call and co-funded by the European Union (H2020, GA No 101017733). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or of the Ministry of University and Research. Neither the European Union nor the granting authority can be held responsible for them. This work was supported by the Provincia Autonoma di Trento, and Q@TN, the joint lab between University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. S.B. acknowledges CINECA for the use of HPC resources under ISCRA-C project DISYK (HP10CGNZG9).
@@
Appendix A Derivation of analytical of magnetization and entanglement for imperfect Bell states
@@
In this appendix, we present the analytical form of the magnetization and quantum entanglement computed for the set of left and right eigenvectors of the non-Hermitian Hamiltonian $H_{\mathrm{nh}}^{\mathrm{Bell}}$. Values of $m_{z}$ evaluated for the ensembles of right and left eigenvectors of $H_{\mathrm{nh}}^{\mathrm{Bell}}$ are given in the following table:
Index ($m$) $m_{z}=\text{Tr}(\hat{\sigma}_{z}\rho_{m}^{L})$, $\rho_{m}^{L}=\text{Tr}_{1}[|L_{m}\rangle\langle L_{m}|]$ $m_{z}=\text{Tr}(\hat{\sigma}_{z}\rho_{m}^{R})$, $\rho_{m}^{R}=\text{Tr}_{1}[|R_{m}\rangle\langle R_{m}|]$
@@ 4 $\frac{1}{2}$ $\frac{1}{2}+\frac{\sqrt{1-\frac{4}{(2+(\alpha-1)^{2})^{2}}}}{2}$
Appendix B Purity score for maximally mixed input state
In this section, we give the analytical derivation for the purity score of the model given by imperfect Bell states. The purity of the ensemble evolved under $H_{\mathrm{nh}}^{\mathrm{Bell}}$ is
$\mathcal{P}(\rho_{RR}(t))=\frac{\text{Tr}\Big{[}\rho_{RR}(t)^{2}\Big{]}}{\text{Tr}(\rho_{RR}(t))^{2}}=\frac{\sum_{mnm^{\prime}n^{\prime}}p_{mn}p_{m^{\prime}n^{\prime}}\langle R_{n^{\prime}}|R_{m}\rangle\langle R_{n}|R_{m^{\prime}}\rangle}{\Big{(}\sum_{mn}p_{mn}\langle R_{n}|R_{m}\rangle\Big{)}^{2}}\,,$ (21)
with $p_{mn}=e^{-i(\lambda_{m}-\lambda_{n})t}\langle L_{m}|\rho_{\mathrm{in}}|L_{n}\rangle$. Similarly, for the quantum state evolved under $(H^{\mathrm{Bell}}_{\mathrm{nh}})^{\dagger}$, we get
$\mathcal{P}(\rho_{LL}(t))=\frac{\text{Tr}\Big{[}\rho_{LL}(t)^{2}\Big{]}}{\text{Tr}(\rho_{LL}(t))^{2}}=\frac{\sum_{mn,m^{\prime}n^{\prime}}{\tilde{p}}_{mn}{\tilde{p}}_{m^{\prime}n^{\prime}}\langle L_{n^{\prime}}|L_{m}\rangle\langle L_{n}|L_{m^{\prime}}\rangle|}{\Big{(}\sum_{mn}\tilde{p}_{mn}\langle L_{n}|L_{m}\rangle\Big{)}^{2}}\,,$ (22)
with $\tilde{p}_{mn}=e^{-i(\lambda_{m}-\lambda_{n})t}\langle R_{m}|\rho_{\mathrm{in}}|R_{n}\rangle$.
Now, for a maximally mixed initial state $\rho=\frac{\mathbb{I}}{4}$, from Eq. (21) we get
$\mathcal{P}(\rho_{RR}(t))$ $=$ $\frac{\sum_{mnm^{\prime}n^{\prime}}\langle L_{m}|L_{n}\rangle\langle L_{m^{\prime}}|L_{n^{\prime}}\rangle\langle R_{n^{\prime}}|R_{m}\rangle\langle R_{n}|R_{m^{\prime}}\rangle}{\Big{(}\sum_{mn}\langle L_{m}|L_{n}\rangle\langle R_{n}|R_{m}\rangle\Big{)}^{2}}$ (23)
[…]

### v2

@@ Acknowledgements.
This project is funded by the European Union - Next Generation EU, Mission 4, Component 2 - CUP E53D23002240006. This project has received funding from the Italian Ministry of University and Research (MUR) through the FARE grant for the project DAVNE (Grant R20PEX7Y3A). The project is funded under the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.4 - Call for tender No. 1031 of 17/06/2022 of Italian Ministry for University and Research funded by the European Union – NextGenerationEU (proj. nr. CN_00000013). This work was supported by the Swiss State Secretariat for Education, Research and innovation (SERI) under contract number UeMO19-5.1. Project DYNAMITE QUANTERA2_00056 funded by the Ministry of University and Research through the ERANET COFUND QuantERA II – 2021 call and co-funded by the European Union (H2020, GA No 101017733). This work was supported by the Caritro Foundation, the Provincia Autonoma di Trento, and Q@TN, the joint lab between University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. S.B. acknowledges CINECA for the use of HPC resources under Italian SuperComputing Resource Allocation– ISCRA Class C Projects No. DISYK-HP10CGNZG9 and DeepSYK - HP10CAD1L3. S.S.R. acknowledges the faculty research scheme at IIT (ISM) Dhanbad, India, under Project No. FRS/2024/PHYSICS MISC0110. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or of the Ministry of University and Research. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Non-Hermiticity quantifiers using other norms and observables
This section provides additional results that reinforce our main conclusions. Beyond the operator norm employed in the definition of $\mathcal{D}$ in the main text, alternative norms–such as the Frobenius norm–can also be considered. When defined with respect to the Frobenius norm, we denote the Hamiltonian non-Hermiticity as $\mathcal{D}^{\prime}$. The numerator of $\mathcal{D}^{\prime}$ is then given by ${\mathcal{F}}=\sqrt{\mathrm{Tr}\Big(\left[\hat{H}^{\mathrm{HN}}_{\mathrm{nh}}-(\hat{H}^{\mathrm{HN}}_{\mathrm{nh}})^{\dagger}\right]^{\dagger}\left[\hat{H}^{\mathrm{HN}}_{\mathrm{nh}}-(\hat{H}^{\mathrm{HN}}_{\mathrm{nh}})^{\dagger}\right]\Big)}$. To compute it explicitly, we start from
$\hat{H}-\hat{H}^{\dagger}=\sum_{n}\left(\varepsilon_{n}|R_{n}\rangle\langle L_{n}|-\varepsilon_{n}^{*}|L_{n}\rangle\langle R_{n}|\right),$ (9)
and use the decompositions
$|R_{n}\rangle$ $=\sum_{m}\alpha_{nm}|R_{m}\rangle+\beta_{nm}|R_{n\perp m}\rangle,\qquad\alpha_{nm}=\langle R_{m}|R_{n}\rangle,\qquad\beta_{nm}=\langle R_{n\perp m}|R_{n}\rangle,$ (10)
$|L_{n}\rangle$ $=\sum_{m}\tilde{\alpha}_{nm}|L_{m}\rangle+\tilde{\beta}_{nm}|L_{n\perp m}\rangle,\qquad\tilde{\alpha}_{nm}=\langle L_{m}|L_{n}\rangle,\qquad\tilde{\beta}_{nm}=\langle L_{n\perp m}|L_{n}\rangle,$ (11)
with the biorthogonality condition $\langle L_{m}|R_{n}\rangle=\delta_{mn}$, and where $|R_{n\perp m}\rangle$ represents the component of $|R\rangle_{n}$ perpendicular to $|R\rangle_{m}$ [11]. The result is
$\mathcal{F}=\sqrt{\text{Tr}(|\hat{H}-\hat{H}^{\dagger}|^{2})}=\sqrt{\sum_{m,n}\left(\varepsilon_{m}^{*}\varepsilon_{n}\alpha_{nm}\tilde{\alpha}_{nm}^{*}+\varepsilon_{n}^{*}\varepsilon_{m}\tilde{\alpha}_{nm}\alpha_{nm}^{*}\right)-\sum_{m}\left(\varepsilon_{m}^{*2}+\varepsilon_{m}^{2}\right)}.$ (12)
As this result shows, the Frobenius norm captures both the influence of non-real energy eigenvalues $\varepsilon_{n}$ as well as non-trivial overlaps $\alpha_{nm}$, $\tilde{\alpha}_{nm}$. Figure 6(a) shows the behavior of $\mathcal{D}^{\prime}$ in the interacting non-Hermitian Hatano–Nelson model. Notably, it exhibits qualitatively similar trends to the operator norm $\mathcal{D}$ analyzed in the main text [see Fig. 1(b)]. It is well conceivable that other relevant functions of $\varepsilon_{n}$, $\alpha_{nm}$, and $\tilde{\alpha}_{nm}$ can be constructed to capture (or even better showcase) the salient features of the system.
As noted earlier, $\mathcal{D}^{\prime}$ becomes nonzero in approximately the same region as the PT breaking transition. In Fig. 7, we present the system-size dependence of $\mathcal{D}^{\prime}$ and compare it with the PT symmetry-breaking transition identified by the nonanalyticity in $\Delta_{01}$, in order to examine in how far both agree in the large-system limit. In Fig. 7(a), we show the variation of $\mathcal{D}^{\prime}$ and its derivative $\mathrm{d}\mathcal{D}^{\prime}/\mathrm{d}V$, while panel (c) shows $\Delta_{01}$, all as a function of $V$ with fixed $\chi=2.5$. The onset of nonanalyticity marks the PT symmetry-breaking transition $\Delta_{01}$. The Hamiltonian non-Hermiticity evolves smoothly across this transition region rather than developing a sharper transition with increasing $N$.
To further quantify the difference between both observables, we derive the critical point $V^{\mathrm{c}}_{\Delta}$ (defined as the peak position of $\Delta_{01}$) and the crossover point $V^{\mathrm{c}}_{\mathcal{D}^{\prime}}$ (given by the minimum of the derivative of $\mathcal{D}^{\prime}$}), whose system-size scaling is shown in panels (b) and (d). The two measures converge to distinct values as $N\to\infty$. This quantitative difference may perhaps be unsurprising: The PT-breaking transition corresponds to an exceptional point between two lowest excited states, whereas the Hamiltonian non-Hermiticity reflects global characteristics of the Hamiltonian across its entire spectrum, as is also evident from Eq. (12).
Furthermore, to extend our analysis of the non-Hermiticity score of observables to non-local observables, we consider the staggered magnetization $\hat{I}=\sum_{i}(-1)^{i}\hat{n}_{i}$. Figure 6(b) shows the trend of $SC^{\hat{I}}_{\rm max}$ across the $\chi$-$V$ parameter space. This non-local observable exhibits behavior qualitatively similar to the site occupation shown in Fig. 4(a).
[…]
Figure 7: (a) Variation of $\mathcal{D}^{\prime}$ and $\frac{\mathrm{d}\mathcal{D}^{\prime}}{\mathrm{d}V}$ with respect to $V$ for fixed $\chi=2.5$. The different curves correspond to system sizes $N=4,6,8,10,12$ and $14$. We consider the minimum of $\frac{\mathrm{d}\mathcal{D}^{\prime}}{\mathrm{d}V}$ to locate the transition point $V_{\mathcal{D}^{\prime}}^{\rm c}$. (b) Finite-size scaling analysis of $V_{\mathcal{D}^{\prime}}^{\rm c}$. The fit approximates $V_{\mathcal{D}^{\prime}}^{\rm c}\approx 0.72\pm 0.02$ at $N\approx\infty$. (c) Variation of $\Delta_{01}$ with $V$ for different system sizes. The non-analytic point in $\Delta_{01}$ locates the critical point $V_{\Delta}$ for the PT transition. (d) Finite size analysis of $\Delta_{01}^{\rm c}$. The fit approximates $\Delta_{01}^{\rm c}\approx 1.72\pm 0.005$ at $N\approx\infty.$
Appendix B Behavior of $\tilde{\mathcal{D}}=|\hat{H}^{\mathrm{HN}}_{\mathrm{nh}}-(\hat{H}^{\mathrm{HN}}_{\mathrm{nh}})^{\dagger}|$ in the $V-\chi$ plane
In this appendix, we present numerical results using—rather than the normalized Hamiltonian non-Hermiticity score, Eq. (1)—the unnormalized norm $\tilde{\mathcal{D}}=|\hat{H}^{\mathrm{HN}}_{\mathrm{nh}}-(\hat{H}^{\mathrm{HN}}_{\mathrm{nh}})^{\dagger}|$. Figure 8 shows the results computed for the interacting Hatano–Nelson model. As may have been expected, the distance between $\hat{H}^{\mathrm{HN}}_{\mathrm{nh}}$ and its Hermitian conjugate $(\hat{H}^{\mathrm{HN}}_{\mathrm{nh}})^{\dagger}$ depends only on the strength of the asymmetric hopping $\chi$. As a consequence, the use of the unnormalized form would suggest physics that remains completely independent of the interaction strength $V$. In particular, it does not capture in any way the changes that occur in the observable scores as $V$ is increased [see Figs. 4 and 5 of the main text]. Moreover, at large interaction the non-Hermitian part becomes insignificant in comparison to the interactions. We thus find it more suitable to normalize by the operator norm $|\hat{H}^{\mathrm{HN}}_{\mathrm{nh}}|$ to obtain a non-Hermitian score $\mathcal{D}$ corresponding to a Hamiltonian with fixed bandwidth [as given in Eq. (1) and Fig. 1(b) of the main text].
[…]
Appendix C Generation of left and right ensembles
$\hat{H}_{\mathrm{nh}}=\sum_{m}\lambda_{m}|R_{m}\rangle\langle L_{m}|,$ (13)
$\hat{\rho}_{RR}(t)$ $=$ $e^{-i\hat{H}_{\mathrm{nh}}t}\hat{\rho}_{\mathrm{in}}e^{it\hat{H}_{\mathrm{nh}}^{\dagger}},$ (14)
where $a_{mn}=\langle L_{m}|\hat{\rho}_{\mathrm{in}}|L_{n}\rangle$. We can distinguish three situations.
Under this condition, Eq. (14) can be rewritten as
$\hat{\rho}_{RR}(t)$ $=\sum_{m}e^{2t\Im[\lambda_{m}]}a_{mm}|R_{m}\rangle\langle R_{m}|$ (15)
At large times, $\hat{\rho}_{RR}(t)$ eventually converges to the right eigenvector with slowest decay rate [5, 36],
[…]
2. All $\lambda_{m}$’s are real (Parity-Time symmetric case).
In this case, Eq. (14) reduces to
$\hat{\rho}_{RR}(t)$ $=\sum_{m}a_{mm}|R_{m}\rangle\langle R_{m}|$
$\hat{\rho}_{RR}(t)\xrightarrow{t\rightarrow\infty}\sum_{m}a_{mm}|R_{m}\rangle\langle R_{m}|.$ (18)
3. Arbitrary non-Hermitian time evolution.
In the general scenario, an observable $\hat{O}$ undergoes the evolution
$\text{Tr}(\hat{O}\hat{\rho}_{RR}(t))$ $=$ $\sum_{mn}e^{-it(\lambda_{m}-\lambda_{n}^{*})}a_{mn}\langle R_{n}|\hat{O}|R_{m}\rangle\,.$ (19)
All the scenarios considered above illustrate the physical origin of expectation values evaluated in the right eigenvectors of the model. A similar analysis with $\hat{H}_{\mathrm{nh}}^{\dagger}$ justifies the same reasoning for the left eigenvectors ($\{\langle L_{m}|\}$). As these considerations also show, while possibly containing interesting properties of the model [36], terms like $\langle R_{m}|\hat{O}|L_{n}\rangle$ do not naturally appear in the evaluation of physical quantities for the quantum state evolved under any generic non-Hermitian dynamics.
@@ Appendix C Generation of left and right ensembles
Appendix D Derivation of analytical of magnetization and entanglement for imperfect Bell states
@@
In this appendix, we present the analytical form of the magnetization and quantum entanglement computed for the set of left and right eigenvectors of the non-Hermitian Hamiltonian $\hat{H}_{\mathrm{nh}}^{\mathrm{Bell}}$. Values of $m_{z}$ evaluated for the ensembles of right and left eigenvectors of $\hat{H}_{\mathrm{nh}}^{\mathrm{Bell}}$ are given in the following table:
Index ($m$) $m_{z}=\text{Tr}(\hat{\sigma}_{z}\hat{\rho}_{m}^{L})$, $\hat{\rho}_{m}^{L}=\text{Tr}_{1}[|L_{m}\rangle\langle L_{m}|]$ $m_{z}=\text{Tr}(\hat{\sigma}_{z}\hat{\rho}_{m}^{R})$, $\hat{\rho}_{m}^{R}=\text{Tr}_{1}[|R_{m}\rangle\langle R_{m}|]$
@@ 4 $\frac{1}{2}$ $\frac{1}{2}+\frac{\sqrt{1-\frac{4}{(2+(\alpha-1)^{2})^{2}}}}{2}$
Appendix E Purity score for maximally mixed input state
In this section, we give the analytical derivation for the purity score of the model given by imperfect Bell states. The purity of the ensemble evolved under $\hat{H}_{\mathrm{nh}}^{\mathrm{Bell}}$ is
$\mathcal{P}(\hat{\rho}_{RR}(t))=\frac{\text{Tr}\Big[\hat{\rho}_{RR}(t)^{2}\Big]}{\text{Tr}(\hat{\rho}_{RR}(t))^{2}}=\frac{\sum_{mnm^{\prime}n^{\prime}}p_{mn}p_{m^{\prime}n^{\prime}}\langle R_{n^{\prime}}|R_{m}\rangle\langle R_{n}|R_{m^{\prime}}\rangle}{\Big(\sum_{mn}p_{mn}\langle R_{n}|R_{m}\rangle\Big)^{2}}\,$, (20)
with $p_{mn}=e^{-i(\lambda_{m}-\lambda_{n})t}\langle L_{m}|\hat{\rho}_{\mathrm{in}}|L_{n}\rangle$. Similarly, for the quantum state evolved under $(\hat{H}^{\mathrm{Bell}}_{\mathrm{nh}})^{\dagger}$, we get
$\mathcal{P}(\hat{\rho}_{LL}(t))=\frac{\text{Tr}\Big[\hat{\rho}_{LL}(t)^{2}\Big]}{\text{Tr}(\hat{\rho}_{LL}(t))^{2}}=\frac{\sum_{mn,m^{\prime}n^{\prime}}{\tilde{p}}_{mn}{\tilde{p}}_{m^{\prime}n^{\prime}}\langle L_{n^{\prime}}|L_{m}\rangle\langle L_{n}|L_{m^{\prime}}\rangle|}{\Big(\sum_{mn}\tilde{p}_{mn}\langle L_{n}|L_{m}\rangle\Big)^{2}}\,$, (21)
with $\tilde{p}_{mn}=e^{-i(\lambda_{m}-\lambda_{n})t}\langle R_{m}|\hat{\rho}_{\mathrm{in}}|R_{n}\rangle$.
Now, for a maximally mixed initial state $\hat{\rho}=\frac{\mathbb{I}}{4}$, from Eq. (20) we get
$\mathcal{P}(\hat{\rho}_{RR}(t))$ $=$ $\frac{\sum_{mnm^{\prime}n^{\prime}}\langle L_{m}|L_{n}\rangle\langle L_{m^{\prime}}|L_{n^{\prime}}\rangle\langle R_{n^{\prime}}|R_{m}\rangle\langle R_{n}|R_{m^{\prime}}\rangle}{\Big(\sum_{mn}\langle L_{m}|L_{n}\rangle\langle R_{n}|R_{m}\rangle\Big)^{2}}$ (22)
$=$ $\frac{\sum_{mnm^{\prime}n^{\prime}}\langle L_{n}|L_{m}\rangle\langle L_{n^{\prime}}|L_{m^{\prime}}\rangle\langle R_{m}|R_{n^{\prime}}\rangle\langle R_{m^{\prime}}|R_{n}\rangle}{\Big(\sum_{mn}\langle L_{n}|L_{m}\rangle\langle R_{m}|R_{n}\rangle\Big)^{2}}\ \text{(taking the conjugate of all the overlaps)}$